# Multi-Level Embedding and Alignment Network with Consistency and Invariance Learning for Cross-View Geo-Localization

### v1

@@ III-E Multi-Loss Optimization
where and , are denoted index of the drone and satellite images, respectively. is the number of images we have used for each view in the training phase. is used to promote local semantic consistency and invariance between and , and defined as follows,
@@
InfoNCE Loss: In the PEE branch, although diversity and hierarchical embedding generation have been achieved, there exists a limitation of the consistency and discriminative capability. To address this issue, we introduce the InfoNCE loss to optimize the embedding space. By contrasting embeddings from different scenes (negative samples) and embeddings from the same scene (positive samples), InfoNCE loss ensures the consistency of embedding features within the same scene while maximizing the discriminability between embeddings of different scenes. The InfoNCE loss contains two queries from both detections, one of which uses a known satellite image to query drone images, namely satellite drone. The other uses a known drone image to query satellite images, which can be termed as drone satellite. The InfoNCE Loss is defined as follows,
where described in Eq.(4) represents an encoded satellite image, which is referred to as the query. is the number of encoded drone images , , called references, among which only one positive sample denoted as matches the query . Similarly, described in Eq.(4) represents an encoded satellite image, which is also called the query. is the number of encoded satellite images , , among which only one positive sample denoted matches the query . The operation represents the sum of the loss for samples during the training phase. is the temperature scaling parameter.
@@ IV-A Experimental Datasets and Evaluation Metrics
University-1652 is a cross-view geo-localization dataset consisting of drone, satellite, and ground-level images from 1,652 locations across 72 universities worldwide. The training set includes 701 buildings from 33 universities, while the test set consists of 951 buildings from 39 universities, with no overlap between training and testing data. This dataset is first used to introduce drone-view to cross-view geo-localization, extended the visual localization task as ground-drone-satellite cross-view matching.
SUES-200 introduces altitude variation in drone images, featuring 200 unique locations. The training and test sets contain 120 and 80 locations, respectively. Each location includes a satellite image and drone images taken at four altitudes: 150m, 200m, 250m, and 300m, encompassing diverse environments such as parks, lakes, and buildings. This dataset evaluates the model’s capability for cross-view retrieval with altitude variations.
Multi-weather University-1652 is an extension of University-1652 with ten simulated weather conditions, including fog, rain, and snow, providing a benchmark to test the model’s robustness under various environmental conditions.
Furthermore, we employ Recall@K (R@K) and Average Precision (AP) as evaluation metrics. R@K measures the proportion of correct matches within the top- retrieved results, while AP represents the balance between precision and recall. Additionally, we evaluate model efficiency through parameter count and computational complexity (GFLOPs) to gauge model portability under resource-constrained conditions.
@@
These datasets present unique challenges such as altitude variation (SUES-200), large-scale distractor sets (University-1652), and weather-induced feature distortion (Multi-weather University-1652), facilitating a comprehensive assessment of cross-view geo-localization performance.
Drone→Satellite | Satellite→Drone | ||||||
@@ IV-B Implementation Details
We adopt a symmetric sampling strategy to select the input images. The ConvNeXt-Tiny model, pre-trained on ImageNet, is used as the backbone network for feature extraction, with a newly added classifier module initialized by the Kaiming initialization method. During both training and testing, all input images are uniformly resized to 3×384×384. We also apply a series of data augmentation techniques, including random cropping, random horizontal flipping and random rotation. The batch size is set to 64 images (32 drone images and 32 satellite images per batch). For optimization, we use the AdamW optimizer with an initial learning rate of 0.001. In , we employ label smoothing with a smoothing factor of 0.1, and the temperature parameter is set as a learnable parameter. Furthermore, in , we introduce two balancing factors, and , to adjust the model’s learning efficacy. All experiments are conducted on the Pytorch deep learning framework, with the experimental platform running on Ubuntu 22.04, equipped with four NVIDIA RTX 4090 GPUs (24GB memory each), leveraging this computationally accelerated environment for experimentation.
@@ IV-C Comparison with State-of-the-art Methods
We compare the proposed MEAN method with several state-of-the-art methods to demonstrate the superiority of our method. The experimental results on the University1652 dataset are reported in Table I, while the results on the Multi-weather University-165 dataset in Table II. Additionally, the results on the SUES-200 dataset are presented in Tables III and IV.
Results on University-1652: As shown in Table I, our MAEN achieves 93.55% R@1 and 94.53% AP in the Drone Satellite setting, and 96.01% R@1 and 92.08% AP in the Satellite Drone setting. These results demonstrate the strong performance and generalization capability of MAEN in cross-view geo-localization. MAEN offers an efficient and lightweight solution, reducing model size by 62.17% and computational complexity by 70.99% compared with the current state-of-the-art model DAC[19], while maintaining comparable performance. Furthermore, with minimal parameter count and computational complexity, MAEN outperforms other advanced methods, including MCCG[7] for multi-feature representation, MFJR[17] for multi-branch joint optimization, and Sample4Geo[5], which employs contrastive optimization techniques. These findings underscore MAEN’s capability to balance high accuracy with computational efficiency, making it a highly effective model for cross-view geo-localization tasks.
@@
MEAN(Ours) | 96.58/89.93 | 96.00/89.49 | 95.15/88.87 | 94.44/87.44 | 93.58/86.91 | 94.44/87.44 | 93.72/86.91 | 96.29/89.87 | 92.87/79.66 | 95.44/86.05 |
Results on Multi-weather University-1652: As shown in Table II, MEAN consistently achieves superior performance across various weather conditions in both Drone Satellite setting and Satellite Drone setting. In the Drone Satellite setting, MEAN achieves the best performance in 8 out of 10 weather conditions. It records R@1 and AP scores of 90.81% and 97.32% under Normal conditions, 90.97% and 92.52% under Fog, and 88.19% and 90.05% under Rain, significantly outperforming other methods. Under challenging conditions like Fog+Rain and Snow, MEAN continues to deliver strong results, achieving an R@1 of 88.60% and AP of 88.84% in Fog+Rain, and 87.21% and 89.21% in Snow. Similarly, in the Satellite Drone setting, MEAN sets new benchmarks in 8 out of 10 weather scenarios. It achieves R@1 scores of 96.58%, 96.00%, and 95.15% under Normal, Fog, and Rain conditions, respectively. Even in adverse scenarios like Over-exposure and Dark, MEAN maintains robust performance, reaching R@1 scores of 92.87% and 96.29%, respectively. Compared with the state-of-the-art models such as Sample4Geo, LPN, and MuSeNet, MEAN demonstrates a remarkable balance between performance and robustness across diverse environmental conditions. These results highlight MEAN’s ability to effectively learn feature consistency and maintain high accuracy, even under challenging cross-view and multi-weather scenarios.
Drone→Satellite | |||||||||||
@@
Results on SUES-200: As shown in Table III, in the Drone Satellite setting, MEAN achieves R@1 scores of 95.50%, 98.38%, 98.95%, and 99.17% and AP scores of 95.46%, 98.72%, 99.05%, and 99.63% at different altitude settings (150m, 200m, 250m, 300m). Although MEAN narrowly falls behind other models at 150m in terms of R@1 and AP, it surpasses state-of-the-art models at the other three altitude levels, achieving the best performance. This demonstrates MEAN’s strong adaptability in higher-altitude domains, effectively preserving high-level semantic consistency and robust across different views. Similarly, as shown in Table IV, in the Satellite Drone setting, MEAN achieves R@1 scores of 97.50%, 97.00%, 98.10%, and 98.87%, and AP scores of 94.75%, 96.06%, 98.06%, and 99.21% across these altitude levels. MEAN consistently achieves the best performance at all heights, further confirming its stability and robustness in extracting and matching cross-view image features, regardless of altitude variations. In general, MEAN demonstrates substantial advantages in both performance and efficiency. Compared with the state-of-the-art models such as Sample4Geo, DAC, and CAMP[30], MEAN achieves a significant reduction in parameter count and computational complexity while maintaining outstanding accuracy. Among 16 evaluation metrics, MEAN achieves the best results in 14 metrics and ranks second in the remaining two, underscoring its robust ability to learn feature consistency and invariance when addressing cross-view perspective differences and scale variations. In addition to its ability to learn consistent and invariant features between cross-view images, MEAN exhibits high computational efficiency, making it an effective solution for cross-view geo-localization in large-scale scenarios.
@@ IV-D Comparison with State-of-the-Art Methods on Cross-Domain Generalization Performance
As shown in Table V and Table VI, MEAN demonstrates exceptional cross-domain adaptability in both Drone Satellite setting and Satellite Drone setting across varying altitudes (150m, 200m, 250m, and 300m). In the Drone Satellite setting (Table V), MEAN achieves competitive results at all altitudes. At 150m, it attains an R@1 of 81.73% and an AP of 87.72%, surpassing other methods. As altitude increases, MEAN maintains robust performance, achieving R@1/AP scores of 93.60%/94.65% at 250m and 94.63%/95.76% at 300m. These results highlight MEAN’s strong generalization capability across varying domains. In the Satellite Drone setting (Table VI), MEAN similarly exhibits superior generalization performance. At 150m, it achieves the highest R@1 of 81.50% and AP of 85.60%, maintaining high competitiveness at higher altitudes. At 250m and 300m, MEAN outperforms other models with R@1/AP scores of 92.06%/93.23% and 92.87%/94.89%, respectively, further validating its robustness in cross-view geo-localization. Compared with the state-of-the-art models such as Sample4Geo, DAC, and CAMP, MEAN achieves superior performance while significantly reducing model complexity. With only 36.50M parameters and 26.18 GFLOPs, MEAN achieves the best performance in 14 of 16 evaluation metrics and secures the second-best results in the remaining two. These findings underscore MEAN’s ability to learn domain-invariant features and adapt effectively to unseen data domains, making it a highly efficient and accurate model for cross-view geo-localization under cross-domain conditions.
@@ IV-E Ablation Studies
In the ablation study, we evaluate the contribution of each component in MEAN on the University-1652 dataset. As shown in Table LABEL:Ablation_Studies, the PEE branch, GEE branch, CEA branch, and CDA Loss are analyzed to assess their individual and combined impacts. Here, and represent the mean squared error and cosine similarity constraints within the CDA Loss, respectively. The baseline model is defined as ConvNeXt-Tiny trained solely with the CE loss, without incorporating any additional branches.
@@
For loss functions, the joint application of and further improves the performance on cross-domain tasks. With all components and losses integrated, MEAN achieves the optimal results: R@1 and AP scores of 93.55% and 94.53% in the Drone Satellite setting, and 96.01% and 92.08% in the Satellite Drone setting. Compared to baseline, MEAN improves R@1 and AP by 9.55% and 7.92% in the Drone Satellite setting, and by 3.79% and 9.28% in the Satellite Drone setting. These results validate the effectiveness of each component in advancing cross-view geo-localization.
@@
Drone→Satellite | Satellite→Drone | |||||||
@@
Feature distribution.To comprehensively evaluate the effectiveness of the MEAN model in cross-view geo-localization, we visualized the intra-class and inter-class distance distributions on the University-1652 dataset, as shown in Fig. 4(a-e). The analysis includes a comparison of two representative methods, MCCG and Sample4Geo, along with a baseline model to validate the contributions of each component in MEAN. Compared with the initial features (Fig. 4(a)), MCCG (Fig. 4(b)), Sample4Geo (Fig. 4(c)), and the baseline model (Fig. 4(d)), the MEAN model achieves substantial separation between intra-class and inter-class distances, resulting in more compact intra-class features and more distinctly separated inter-class features. However, both the baseline model and the MCCG exhibit limitations in intra-class compactness, inter-class separation, and cross-view consistency. Although Sample4Geo demonstrates improved discriminative ability and consistency over the baseline and MCCG, it still exhibits overlap in some areas, indicating suboptimal intra-class compactness and inter-class separation. By contrast, the MEAN model (Fig. 4(e)) significantly reduces intra-class distances while increasing inter-class distances, showcasing superior feature discriminability and cross-view consistency.
Fig.4(f-j) further illustrates the feature distributions of different models in a 2D feature space obtained using t-SNE[39]. The initial features (Fig.4(f)) and the baseline model (Fig.4(i)) show poor separation between features from different locations, whereas features from the same location are widely dispersed. In comparison, MCCG (Fig.4(g)) and Sample4Geo (Fig.4(h)) show some improvements in intra-class compactness and inter-class separation, though partial overlap persists between features of different locations. In contrast, the MEAN model (Fig.4(j)) demonstrates a distinct advantage: images from the same location, including UAV and satellite views, are closely clustered in feature space, while features from different locations are clearly separated. This distribution highlights the strong modality invariance of the MEAN model, effectively mitigating feature discrepancies caused by viewpoint variations and enhancing cross-view consistency.
IV-F Visualization
Retrieval result. To further illustrate the effectiveness of the MEAN model, we present the retrieval results on the University-1652 dataset, as shown in Fig. 5. For each retrieval result, the green border indicates correctly matched images, while the red border signifies incorrect matches. The results indicate that the MEAN model significantly improves retrieval performance.

### v2

@@ III-E Multi-Loss Optimization
where and , are denoted index of the aerial and satellite images, respectively. is the number of images we have used for each view in the training phase. is used to promote local semantic consistency and invariance between and , and defined as follows,
@@
InfoNCE Loss: In the PEE branch, although diversity and hierarchical embedding generation have been achieved, there exists a limitation of the consistency and discriminative capability. To address this issue, we introduce the InfoNCE loss to optimize the embedding space. By contrasting embeddings from different scenes (negative samples) and embeddings from the same scene (positive samples), InfoNCE loss ensures the consistency of embedding features within the same scene while maximizing the discriminability between embeddings of different scenes. The InfoNCE loss contains two queries from both detections, one of which uses a known satellite image to query aerial images, namely satellite drone. The other uses a known aerial image to query satellite images, which can be termed as dronesatellite. The InfoNCE Loss is defined as follows,
where described in Eq.(4) represents an encoded satellite image, which is referred to as the query. is the number of encoded aerial images , , called references, among which only one positive sample denoted as matches the query . Similarly, described in Eq.(4) represents an encoded satellite image, which is also called the query. is the number of encoded satellite images , , among which only one positive sample denoted matches the query . The operation represents the sum of the loss for samples during the training phase. is the temperature scaling parameter.
@@ IV-A Experimental Datasets and Evaluation Metrics
University-1652 is a cross-view geo-localization dataset consisting of drone, satellite, and ground-level images from 1,652 locations across 72 universities worldwide. The training set includes 701 buildings from 33 universities, while the test set consists of 951 buildings from 39 universities, with no overlap between training and testing data. This dataset is first used to introduce drone view to cross-view geo-localization, extended the visual localization task as ground-drone-satellite cross-view matching.
SUES-200 introduces altitude variation in aerial images, featuring 200 unique locations. The training and test sets contain 120 and 80 locations, respectively. Each location includes a satellite image and aerial images taken at four altitudes: 150m, 200m, 250m, and 300m, encompassing diverse environments such as parks, lakes, and buildings. This dataset evaluates the model’s capability for cross-view retrieval with altitude variations.
Multi-weather University-1652 is an extension of University-1652 with ten simulated weather conditions, including fog, rain, and snow. It provides a benchmark to test the model’s robustness under various environmental conditions.
Furthermore, we employ Recall@K (R@K) and Average Precision (AP) as evaluation metrics. R@K measures the proportion of correct matches within the Top- retrieved results, while AP represents the balance between precision and recall. Additionally, we evaluate model efficiency through parameter count and computational complexity (GFLOPs) to reflect model portability under resource-constrained conditions. In the evaluation of parameter counts and model complexity, comparisons are conducted using the model states that achieve the optimal performance for each method.
@@
These datasets present unique challenges such as altitude variation (SUES-200), large-scale distractor sets (University-1652), and weather-induced feature distortion (Multi-weather University-1652) to facilitate a comprehensive assessment of cross-view geo-localization performance.
| DroneSatellite | SatelliteDrone | ||||||
@@ IV-B Implementation Details
We adopt a symmetric sampling strategy to select the input images. The ConvNeXt-Tiny model, pre-trained on ImageNet, is used as the backbone network for feature extraction with a newly added classifier module initialized by the Kaiming initialization method. During both training and testing, all input images are uniformly resized to . We also apply a series of data augmentation techniques, including random cropping, random horizontal flipping and random rotation. The batch size is set to 64 images (32 aerial images and 32 satellite images per batch). For optimization, we use the AdamW optimizer with an initial learning rate of 0.001. In , we employ label smoothing with a smoothing factor of 0.1, and the temperature parameter is set as a learnable parameter. Furthermore, we introduce two balancing factors and to adjust the model’s learning efficacy in . All experiments are conducted on the Pytorch deep learning framework with the experimental platform running on Ubuntu 22.04 equipped with four NVIDIA RTX 4090 GPUs.
@@ IV-C Comparison with State-of-the-art Methods
We compare the proposed MEAN method with several state-of-the-art methods to demonstrate the superiority of our method. The experimental results on the University1652 dataset are reported in Table I, while the results on the Multi-weather University-165 dataset in Table II. Additionally, the results pertaining to the SUES-200 dataset are shown in Tables III and IV.
Results on University-1652: As shown in Table I, MAEN achieves 93.55% R@1 and 94.53% AP in the DroneSatellite setting, and 96.01% R@1 and 92.08% AP in the SatelliteDrone setting. These results demonstrate the strong performance and generalization capability of MAEN in cross-view geo-localization. MAEN offers an efficient and lightweight solution, reducing model size by 62.17% and computational complexity by 70. 99% compared to the current state-of-the-art model DAC [19], while maintaining comparable performance. Furthermore, with minimal parameter count and computational complexity, MAEN outperforms other advanced methods, including MCCG [7] for multi-feature representation, MFJR [17] for multi-branch joint optimization, and Sample4Geo [5], which employed the contrastive optimization techniques. These findings highlight MAEN’s ability to achieve high accuracy while maintaining computational efficiency, making it an effective model for cross-view geo-localization.
@@
| MEAN(Ours) | 96.58/89.93 | 96.00/89.49 | 95.15/88.87 | 94.44/87.44 | 93.58/86.91 | 94.44/87.44 | 93.72/86.91 | 96.29/89.87 | 92.87/79.66 | 95.44/86.05 | ||
Results on Multi-weather University-1652: As shown in Table II, MEAN consistently achieves superior performance across various weather conditions in both DroneSatellite setting and SatelliteDrone setting. In the DroneSatellite setting, MEAN achieves the best performance across all 10 weather conditions. It records R@1 and AP scores of 90.81% and 92.32% under Normal conditions, 90.97% and 92.52% under Fog, and 88.19% and 90.05% under Rain, significantly outperforming other methods. Under challenging conditions like Snow and Fog+Rain, MEAN continues to deliver strong results, achieving an R@1 of 88.69% and AP of 90.49% in Snow, and 86.75% and 88.84% in Fog+Rain.Results under other conditions show similar trends. Similarly, in the SatelliteDrone setting, MEAN sets new benchmarks in 8 out of 10 weather scenarios. It achieves R@1 scores of 96.58%, 96.00%, and 95.15% under Normal, Fog, and Rain conditions, respectively. Even in adverse scenarios such as Over-exposure and Dark, MEAN maintains robust performance, reaching R@1 scores of 92.87% and 96.29%, respectively. Results under other conditions show similar trends. Compared with the state-of-the-art models such as Sample4Geo, LPN [13], and MuSeNet [36], MEAN demonstrates a remarkable balance between performance and robustness across diverse environmental conditions. These results highlight MEAN’s ability to effectively learn feature consistency and maintain high accuracy, even under challenging cross-view and multi-weather scenarios.
| DroneSatellite | |||||||||||
@@
Results on SUES-200: As shown in Table III, in the DroneSatellite setting, MEAN achieves R@1 scores of 95.50%, 98.38%, 98.95%, and 99.52% and AP scores of 96.46%, 98.72%, 99.17%, and 99.63% at different altitude settings (150m, 200m, 250m, 300m). Although MEAN narrowly falls behind other models at 150m in terms of R@1 and AP, it surpasses state-of-the-art models at the other three altitude levels, achieving the best performance. This demonstrates MEAN’s strong adaptability in higher-altitude domains, effectively preserving high-level semantic consistency and robust across different views. Similarly, as shown in Table IV, in the SatelliteDrone setting, MEAN achieves R@1 scores of 97.50%, 100.00%, 100.00%, and 100.00%, and AP scores of 94.75%, 97.09%, 98.28%, and 99.21% across these altitude levels. MEAN consistently achieves the best performance at all heights, further confirming its stability and robustness in extracting and matching cross-view image features, regardless of altitude variations. In general, MEAN demonstrates substantial advantages in both performance and efficiency. Compared with the state-of-the-art models such as Sample4Geo, DAC, and CAMP [30], MEAN achieves a significant reduction in parameter count and computational complexity while maintaining outstanding accuracy. Among 18 evaluation metrics, MEAN achieves the best results in 16 metrics and ranks second in the remaining two, highlighting its robust ability to learn feature consistency and invariance when addressing cross-view perspective differences and scale variations. In addition to its ability to learn consistent and invariant features between cross-view images, MEAN exhibits high computational efficiency, making it an effective solution for cross-view geo-localization in large-scale scenarios.
@@ IV-D Comparison with State-of-the-Art Methods on Cross-Domain Generalization Performance
As shown in Table V and Table VI, MEAN demonstrates exceptional cross-domain adaptability in both DroneSatellite setting and SatelliteDrone setting across varying altitudes (150m, 200m, 250m, and 300m). In the DroneSatellite setting (Table V), MEAN achieves competitive results at all altitudes. At 150m, it attains an R@1 of 81.73% and an AP of 87.72%, surpassing other methods. As altitude increases, MEAN maintains robust performance, achieving R@1/AP scores of 89.05%/91.00% at 200m, 92.13%/93.60% at 250m and 94.63%/95.76% at 300m. These results highlight MEAN’s strong generalization capability across varying domains. In the SatelliteDrone setting (Table VI), MEAN similarly exhibits superior generalization performance. At 150m, it achieves the highest R@1 of 91.25% and AP of 81.50%, maintaining high competitiveness at higher altitudes. At 200m, 250m and 300m, MEAN outperforms other models with R@1/AP scores of 96.25%/89.55, 95.00%/92.36% and 96.25%/94.32%, respectively, further validating its robustness in cross-view geo-localization. Compared with the state-of-the-art models such as Sample4Geo, DAC, and CAMP, MEAN achieves superior performance while significantly reducing model complexity. With only 36.50M parameters and 26.18 GFLOPs, MEAN achieves the best performance in 13 out of 18 evaluation metrics and ranks second in four metrics. These findings underscore MEAN’s ability to learn domain-invariant features and adapt effectively to unseen data domains, making it a highly efficient and accurate model for cross-view geo-localization under cross-domain conditions.
@@ IV-E Ablation Studies
In the ablation study, we evaluate the contribution of each component in MEAN on the University-1652 dataset. As shown in Table LABEL:Ablation_Studies, the PEE branch, GEE branch, CEA branch, and CDA Loss are analyzed to assess their individual and combined impacts. Here, and represent the mean squared error and cosine similarity constraints within the CDA Loss, respectively. The baseline model is defined as ConvNeXt-Tiny with the CE loss trained solely without incorporating any additional branches.
@@
For loss functions, the joint application of and further improves the performance on cross-domain tasks. With all components and losses integrated, MEAN achieves the optimal results: R@1 and AP scores of 93.55% and 94.53% in the DroneSatellite setting, and 96.01% and 92.08% in the SatelliteDrone setting. Compared to the baseline, MEAN improves R@1 and AP by 9.55% and 7.92% in the DroneSatellite setting, and by 3.79% and 9.28% in the SatelliteDrone setting. These results validate the effectiveness of each component in advancing cross-view geo-localization.
@@
| DroneSatellite | SatelliteDrone | |||||||
@@
IV-F Feature Distribution
To comprehensively evaluate the effectiveness of the MEAN model in cross-view geo-localization, we visualized the intra-class and inter-class distance distributions on the University-1652 dataset, as shown in Fig. 4(a-e). The analysis includes a comparison of two representative methods, MCCG and Sample4Geo, along with a baseline model to validate the contributions of each component in MEAN. Compared with the initial features (Fig. 4(a)), MCCG (Fig. 4(b)), Sample4Geo (Fig. 4(c)), and the baseline model (Fig. 4(d)), the MEAN model achieves substantial separation between intra-class and inter-class distances, resulting in more compact intra-class features and more distinctly separated inter-class features. However, both the baseline model and the MCCG exhibit limitations in intra-class compactness, inter-class separation, and cross-view consistency. Although Sample4Geo demonstrates improved discriminative ability and consistency over the baseline and MCCG, it still exhibits overlap in some areas, indicating suboptimal intra-class compactness and inter-class separation. By contrast, the MEAN model (Fig. 4(e)) significantly reduces intra-class distances while increasing inter-class distances, showcasing superior feature discriminability and cross-view consistency.
Fig.4(f-j) further illustrates the feature distributions of different models in a 2D feature space obtained using t-SNE[39]. The initial features (Fig.4(f)) and the baseline model (Fig.4(i)) show poor separation between features from different locations, whereas features from the same location are widely dispersed. In comparison, MCCG (Fig.4(g)) and Sample4Geo (Fig.4(h)) show some improvements in intra-class compactness and inter-class separation though partial overlap persists between features of different locations. In contrast, the MEAN model (Fig.4(j)) demonstrates a distinct advantage: images from the same location, including drone and satellite views, are closely clustered in feature space, while features from different locations are clearly separated. This distribution highlights the strong modality invariance of the MEAN model, effectively mitigating feature discrepancies caused by viewpoint variations and enhancing cross-view consistency.
IV-G Retrieval Results
To further illustrate the effectiveness of the MEAN model, we present the retrieval results on the University-1652 dataset as shown in Fig. 5. For each retrieval result, the green border indicates correctly matched images, while the red border signifies incorrect matches. The results indicate that the MEAN model significantly improves retrieval performance.